OpenStudy (anonymous):

Solve for x. 7x^2-2x+2=5x^2+x

OpenStudy (anonymous):

I believe you would use the quadratic formula but, I'm not quite sure what to do with the five.

OpenStudy (babalooo):

collect like terms, then use quadratic formula

OpenStudy (anonymous):

Okay. So, I put it all on the right and then, simplified. I got 12x^2-1x+2=0

OpenStudy (babalooo):

are you sure it's 12x^2

OpenStudy (babalooo):

are you sure it's -1x?

OpenStudy (anonymous):

Okay. I'm not sure to be honest. Instead, would 2x^2-3x+2=0 make more sense? I think so.

OpenStudy (anonymous):

I wasn't sure whether to subtract the 5x^2 and 'x' or to just add it onto the right.

OpenStudy (anonymous):

From here may I use the quadratic formula?

OpenStudy (babalooo):

yes, when you pull values from one side to the other, you have to change the sign like 1+1 =2 if you pull the 1 over 1=2+1 does not make sense . 1=2-1

OpenStudy (babalooo):

yes you can use the quadratic equation here if you would like

OpenStudy (anonymous):

Oh okay. :D

OpenStudy (aum):

7x^2-2x+2=5x^2+x Add -5x^2 to both sides: 2x^2 - 2x + 2 = x Add -x to both sides: 2x^2 - 3x + 2 = 0 The discriminant is: b^2 - 4ac = (-3)^2 - 4(2)(2) = 9 - 16 = -7. Since the discriminant is negative, the roots will be complex. plug it into the quadratic formula.

OpenStudy (anonymous):

I got it! :D Thank you babalooo and aum!

OpenStudy (aum):

you are welcome.

OpenStudy (aum):

Discriminant = -7 Square root of discriminant = \(\sqrt{7}i\) \(\Large x = \frac{-b \pm \sqrt{discriminant}}{2a} = \frac{3 \pm \sqrt{7}i}{4}\)